### v1

# Lifshitz Quantum Mechanics and Anisotropic Josephson Junction

Abstract
We consider quantum mechanics in spacetime with time anisotropy. Such Lifshitz quantum mechanics is characterized by a kinetic term with fractional derivatives. We show that, contrary to popular belief, local conservation of probability is respected when the probability current is properly identified. As an application we consider a Josephson Junction with an anisotropic insulating layer. We show that anisotropy affects the tunneling rate and can greatly enhance the working of Josephson Junction.
I Introduction
Many interesting systems in physics are anisotropic. For example, liquid crystal, magnetic materials and critical systems in condensed matter physics are ubiquitously anisotropic. The cosmic microwave background also display traces of anisotropy that are thought to have originated from quantum fluctuations occurred during inflation. Often the fundamental physical law is isotropic but the presence of a background leads to anisotropy. An interesting instant of anisotropy is when physics is invariant under the Lifshitz scaling transformation
$t\rightarrow\lambda^{z}t,\quad x^{i}\rightarrow\lambda x^{i},\quad\lambda>0.$ (I.1)
Here spatial isotropy is maintained while time is anisotropic for an anisotropy exponent $z\neq 1$. Such Lifshitz scaling symmetry occurs naturally as symmetry of the critical point of renormalization group (RG) flow for anisotropic systems. Lifshitz scale invariant system can be recognized easily by the dispersion relation of its massless modes:
$\omega\sim k^{z},$ (I.2)
where one can extract not only the anisotropy exponent $z$ but also some mass scale that is associated with the anisotropy. As the only possible form that is compatible with the scaling symmetry (I.1), (I.2) is independent of the underlying origin of the anisotropy. The symmetry for $z=2$ was first proposed by Lifshitz [1] to describe the tricritical point of liquid helium [2]. The dispersion relation (I.2) is also featured in modified gravity scenario with a ghost condensate [3]. In condensed matter physics, a well known example is the Rokhsar-Kivelson (RK) Quantum dimer model [4, 5] where a continuum $z=2$ Lifshitz scalar field theory emerges at the critical point.
The $z=2$ anisotropic scaling symmetry (I.1) has been intensively studied from the field theory point of view [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 16, 18, 19]. For example, it is known that Lifshitz scalar field theory admits ground state called the RK vacuum, with the properties that it is local and is given by a superposition of quantum states determined by some classical probability distribution [4]. The point $z=2$ is somewhat special since the Lifshitz scaling transformation (I.1) combines with the Galilean boost to form the Schrodinger group only for this value of $z$. Study of field theory with general $z$ was initiated in [20] where with the help of fractional calculus. It was shown that, independent of whether there is an enlarged Schrodinger symmetry or not, the ground state of the Lifshitz scalar field always takes the form of RK. The entropic properties of the RK vacuum were also studied, both in field theory and by holography, in [20].
Clearly, the physics of anisotropy is a subject of its own interest that goes beyond quantum field theory. With general applicability in mind, we explore in this paper the physical effect of anisotropy in quantum mechanics. In particular we consider a kind of time anisotropy of the background space time specified by (I.1). With the dispersion relation fixed to be of the form (I.1), quantum mechanics of free particle in the anisotropic spacetime reads
$i\hbar\frac{\partial}{\partial t}\psi=H_{0}\psi,\quad H_{0}=-\frac{\hbar^{z}}{M^{z-2}}\frac{\Delta_{z}}{2m}$ (I.3)
where $H_{0}$ is the free Hamiltonian, $m$ is the mass of the particle and $M$ is a mass scale associated with the anisotropy. The fractional Laplacian $\Delta_{z}$ of order $z$ is defined by its action on the basis of exponential functions by
$\triangle_{z}\,e^{{\bf k}\cdot{\bf x}}=({\bf k}^{2})^{z/2}e^{{\bf k}\cdot{\bf x}}.$ (I.4)
The free (I.3) is Lifshitz scale invariant. One may also add a potential and consider the Lifshitz quantum mechanics defined by
$i\hbar\frac{\partial}{\partial t}\psi=H\psi,\quad H=-\frac{\hbar^{z}}{M^{z-2}}\frac{\Delta_{z}}{2m}+V.$ (I.5)
Note that in general the scaling symmetry is broken except for some special form of the potential, e.g. $V\sim 1/|x|^{z}$, in which case we obtain the Lifshitz invariant quantum mechanics. In analogous to conformal quantum mechanics [21], it may be interesting to study further this kind of Lifshitz invariant quantum mechanics.
Here is a summary of the paper. In section 2, we consider the Lifshitz quantum mechanics in general. There is a belief in the literature that local conservation of probability is violated by a source term in Lifshitz quantum mechanics. It turns out it is in fact the probability current that is modified by anisotropy. And once the correct probability current is employed, the local conservation of probability is respected. In section 3, we consider an anisotropic Josephson Junction that is made up of an anisotropic insulating layer sandwiched by two normal superconducting materials. We show that the critical current can be greatly enhanced by anisotropy. We argue that anisotropic Josephson Junction has great potential in quantum computing.
II Lifshitz Quantum Mechanics
Consider Lifshitz quantum mechanics (I.5). This kind of quantum mechanics is also refereed to as fractional quantum mechanics, see for example [22] for some discussion. It is easy to show that the Lifshitz quantum mechanics (I.5) is unitary and admits the local conservation law
$\frac{\partial\rho}{\partial t}+\nabla\cdot\pmb{J}=0,$ (II.1)
where the probability density takes the usual form $\rho=\psi^{*}\psi$, while the probability current density is modified by anisotropy. As the fractional derivative is defined using the basis of exponential functions, $\pmb{J}$ can be most easily obtained in the momentum space and we have
$\pmb{J}=-\frac{1}{2mM^{z-2}}\int\frac{d^{D}{p}\,d^{D}{q}}{(2\pi\hbar)^{D}}\,e^{i({\bf p}-{\bf q})\cdot{\bf x}/\hbar}\varphi^{*}({\bf q},t)\varphi({\bf p},t)\frac{(-{\bf q}^{2})^{z/2}-(-{\bf p}^{2})^{z/2}}{{\bf q}^{2}-{\bf p}^{2}}({\bf q}+{\bf p}),$ (II.2)
where $\varphi({\bf p},t)$ is the momentum space wave function defined by the Fourier transform
$\psi({\bf x},t)=\int\frac{d^{D}p}{(2\pi\hbar)^{D/2}}\,e^{i{\bf p}\cdot{\bf x}/\hbar}\varphi({\bf p},t).$ (II.3)
This result can be obtained from the Noether theorem associated with the phase symmetry $\psi\to\psi e^{i\theta}$ of quantum mechanics or by the direct manipulation of the wave equation (I.5). Equivalently in coordinate space, it is
$\mathbf{J}=\lambda_{M}^{z-2}\frac{\hbar}{2mi}\frac{\nabla_{x}-\nabla_{x^{\prime}}}{\nabla_{x}^{2}-\nabla_{x^{\prime}}^{2}}\left[(\nabla_{x}^{2})^{z/2}-(\nabla_{x^{\prime}}^{2})^{z/2}\right]\psi^{*}(\mathbf{x}^{\prime},t)\psi(\mathbf{x},t)\Big{|}_{\mathbf{x}^{\prime}=\mathbf{x}},$ (II.4)
where $\lambda_{M}:=\hbar/M$ is a length scale associated with the anisotropy.
We remark that in the form (II.2), it is easy to see that instead of using the kernel $\frac{{\bf q}+{\bf p}}{{\bf q}^{2}-{\bf p}^{2}}$ in (II.2), one may also use any vector $\bm{\beta}$ that satisfies
$\bm{\beta}\cdot({\bf q}-{\bf p})=1$ (II.5)
to produce a $\pmb{J}$ that observes the conservation law. For example, $\bm{\beta}=\frac{{\bf q}-{\bf p}}{|{\bf q}-{\bf p}|^{2}}$ is good too. To see this, suppose $\bm{\beta}_{2},\bm{\beta}_{1}$ both satisfy (II.5), then $(\bm{\beta}_{2}-\bm{\beta}_{1})$ is perpendicular to $({\bf q}-{\bf p})$, i.e. $\bm{\beta}_{2}-\bm{\beta}_{1}=({\bf q}-{\bf p})\times{\bf s}$ for some ${\bf s}$. This implies that the corresponding $\pmb{J}$’s differ from each other by a curl term, and hence both satisfy the same conservation law. This kind of ambiguity is always there even for the standard quantum mechanics although it is not usually mentioned. Like the Poynting vector, additional input from the physical setting such as symmetry, boundary condition or a good understanding of the underlying dynamics, is usually needed to fix the choice of $\pmb{J}$ properly. This is the case for the example that we will consider below. We remark also that the local conservation of probability in fractional quantum mechanics has been discussed in [23, 24, 25] where a source term has been found, and which has led to the confusion of non-conservation of local probability. As we showed above, the probability current is indeed modified by anisotropy and local conservation of probability is observed in fractional quantum mechanics. This conclusion is important as we will use below the local conservation of probability to derive a statement about the macroscopic electric current in the Josephson Junction system.
III Anisotropic Josephson Junction
As an example of how anisotropy may lead to interesting quantum physics, we consider tunneling through an anisotropic region. A prominent example where the tunneling effect has a macroscopic manifestation is the Josephson effect where cooper pairs of electrons can flow between two superconductors separated by a thin insulating barrier.
Consider a SIS type Josephson Junction that is made up of two superconductors separated by a thin insulator. The two superconductors are assumed to be made of ordinary isotropic materials while the insulator is made of an anisotropic material. The latter can be achieved by material engineering such that the dispersion relation (I.2) arises and the insulator behaves effectively as an anisotropic material. In quantum mechanics, the superconductors have Landau-Ginzburg order parameters
$\psi_{i}=\sqrt{n_{i}}e^{{\rm i}\theta_{i}},\quad i=1,2$ (III.1)
that can be interpreted as the wave functions of the Cooper pairs in the two superconducting regions. See figure 1.
Figure 1: Two superconductors separated by a thin insulator
Suppose the Cooper pairs have typical energy $E$ in the regions 1, 2 and the insulator is represented as a potential barrier $V_{0}>E$ (see figure 2). Then the wave equation inside the insulator,
$\lambda_{M}^{z-2}\frac{\hbar^{2}\Delta_{z}\psi}{2m}=(V_{0}-E)\psi$ (III.2)
has solution
$\psi=C_{1}\cosh\frac{x}{\zeta}+C_{2}\sinh\frac{x}{\zeta},$ (III.3)
where $C_{1},C_{2}$ are arbitrary constants and
$\zeta:=\zeta_{0}\left(\frac{\zeta_{0}}{\lambda_{M}}\right)^{2/z-1},\quad\mbox{with}\;\;\zeta_{0}:=\frac{\hbar}{\sqrt{2m(V_{0}-E)}}.$ (III.4)
Here $\zeta$ is the tunneling penetration length and $\zeta_{0}$ is the tunneling penetration length in the undeformed $z=2$ (isotropic) case. Note that the penetration length is modified by anisotropy through the presence of the anisotropy length scale $\lambda_{M}$ and anisotropy index $z$. Next let us determine the tunneling current.
Figure 2: Tunneling through a potential
Note that here we have assumed the Junction to be uniform in the transverse directions and so we have effectively a one spatial dimension system. In this case there is no room for any curl term and the current is uniquely defined by (II.4). We obtain
$J=\frac{z\zeta}{2\zeta_{0}}\cdot\frac{\hbar}{m\zeta_{0}}{\rm Im}(C_{1}^{*}C_{2}).$ (III.5)
The coefficients $C_{1},C_{2}$ can be determined by imposing the boundary conditions at $x=\pm a$:
$\psi(-a)=\sqrt{n_{1}}e^{{\rm i}\theta_{1}},\quad\psi(a)=\sqrt{n_{2}}e^{{\rm i}\theta_{2}}.$ (III.6)
Taking into account the Cooper pair has charge $q=2e$, we obtain the first Josephson equation for the tunneling current
$J=J_{c}\sin\varphi,\quad\varphi=\theta_{2}-\theta_{1}.$ (III.7)
Here the critical current $J_{c}$ is given by
$J_{c}=\gamma J_{c0}$ (III.8)
with
$J_{c0}=\frac{q\hbar\sqrt{n_{1}n_{2}}}{m\zeta_{0}}\frac{1}{\sinh 2a/\zeta_{0}},\quad\gamma=\frac{z\zeta}{2\zeta_{0}}\frac{\sinh 2a/\zeta_{0}}{\sinh 2a/\zeta},$ (III.9)
where $J_{c0}$ is the critical current in the isotropic case and $\gamma$ is an modification factor due to anisotropy. If we consider the application of a voltage $V$ across the Josephson Junction, then the phase becomes time dependent and we have the second Josephson equation,
$\frac{\partial\varphi}{\partial t}=\frac{q}{\hbar}V.$ (III.10)
Note that this result is unmodified by anisotropy. The derivation is the same as usual. In fact, due to gauge invariance, the wavefunction in the superconductor is modified to have a phase factor
[…]
when an electromagnetic field is introduced. This results in a modification to the order parameter
[…]
where the subscript “0” denotes the quantity before the electric field is introduced. This leads immediately to (III.10).
Note that in the usual isotropic case, the critical current $J_{c0}$ is exponentially suppressed by a factor of $e^{-2a/\zeta}$ since the Cooper pairs cannot tunnel deep enough into the insulating region. However this may change in the anisotropic situation. In fact, the Josephson current can be enhanced ($\gamma>1$) if the penetration length is made bigger ($\zeta>\zeta_{0}$). This happens if $z<2$ and $\lambda_{M}<\zeta_{0}$ or $z>2$ and $\lambda_{M}>\zeta_{0}$. Thus an anisotropic material has the potential of enhancing the working of Josephson Junction. In particular, if the anisotropic insulating material can be engineered such that the penetration length $\zeta$ get to become of the same order as $a$, then tunneling can become much more effective and $J_{c}$ may be greatly enhanced. As illustration, let us consider an isotropic sample with $\zeta_{0}/2a=1/10$, in this case the critical current is suppressed with a factor of $\sim e^{-10}\sim 10^{-5}$. Introducing anisotropy, for example with $z=0.8$ and $\zeta_{0}/\lambda_{M}=10$, we can have a huge amplification factor $\gamma\sim 10^{5}$; or with $z=5$ and $\zeta_{0}/\lambda_{M}=1/10$, we have $\gamma\sim 10^{4}$. We remark that the dispersion relation (I.2) is acausal for $z<1$ due to the existence of superluminal modes [26] and the violation of the null energy conditions of the holographic dual [27]. Therefore a junction with $z\geq 1$ may be more physical. It is interesting to verify this prediction of enhanced tunneling experimentally.
The amplification of critical current due to time anisotropy may be useful for improving the performance of Josephson Junction in quantum computing. Josephson Junction also found useful application in digital circuits as switching elements due to it’s high switching speed and low power dissipation [28]. Some characteristic time constants in Josephson Junction depends on the critical current, an anisotropic Junction can have time constants that are more useful for the switching purpose. Given the wide variety of applications of Josephson Junction [29], anisotropic Josephson Junction appears to be an interesting topic that deserves further study.
Acknowledgments
CSC would like to thank Harold Steinacker for discussions and the physics department of the University of Vienna for hospitality where part of the work was done. The support of this work by NCTS and the grant 113-2112-M-007-039-MY3 of the National Science and Technology Council of Taiwan is gratefully acknowledged.