### v2

# Lifshitz Quantum Mechanics and Anisotropic Josephson Junction

Abstract
We consider quantum mechanics in spacetime with anisotropy in time. Such Lifshitz quantum mechanics is characterized by a kinetic term with fractional derivatives. We show that, contrary to a common claim in the literature, local conservation of probability is respected when the probability current is properly identified. As an application we consider a Josephson Junction with an insulating layer exhibiting Lifshitz anisotropy. We show that anisotropy modifies the tunneling rate and can significantly enhance the performance of the Josephson Junction.
I Introduction
Many interesting systems in physics are anisotropic. For example, liquid crystals, magnetic materials and critical systems in condensed matter physics are ubiquitously anisotropic. The cosmic microwave background also display traces of anisotropy that are thought to have originated from quantum fluctuations occurred during inflation. Often the fundamental physical laws are isotropic but the background structure induces anisotropy. An interesting instance of anisotropy is when physics is invariant under the Lifshitz scaling transformation
$t\rightarrow\lambda^{z}t,\quad x^{i}\rightarrow\lambda x^{i},\quad\lambda>0.$ (I.1)
Here spatial isotropy is maintained while time is anisotropic for a dynamical exponent $z\neq 1$. Such Lifshitz scaling symmetry occurs naturally as the symmetry at the critical point of renormalization group (RG) flow for anisotropic systems. Lifshitz scale invariant system can be recognized easily by the dispersion relation of their massless modes:
$\omega\sim k^{z},$ (I.2)
from which one can extract not only the dynamical exponent $z$ but also a mass scale associated with the anisotropy. As the only possible form that is compatible with the scaling symmetry (I.1), (I.2) is independent of the underlying origin of the anisotropy. The symmetry for $z=2$ was first proposed by Lifshitz [1] to describe the tricritical point of liquid helium [2]. The dispersion relation (I.2) also appears in certain modified gravity scenarios involving ghost condensate [3]. In condensed matter physics, a well known example is the Rokhsar-Kivelson (RK) Quantum dimer model [4, 5] where a continuum $z=2$ Lifshitz scalar field theory emerges at the critical point.
Explicit field theoretic realization of the Lifshitz symmetry (I.1) for the special integer value of $z=2$ has been known as the quantum Lifshitz model (QLM) [6] . Various bipartite entanglement measures were subsequently analyzed within the QLM [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 16, 18, 19]. It is known that Lifshitz invariant scalar field theory admits ground state called the Rokhsar-Kivelson (RK) vacuum [4, 5], with the properties that it is local and is given by a superposition of quantum states determined by some classical probability distribution.
The point $z=2$ is however somewhat special since the Lifshitz scaling transformation (I.1) combines with the Galilean boost to form the Schrodinger group only for this value of $z$. The study of field theory beyond $z=2$ was pioneered in [20] where QLM was generalized to $(D+1)$-dimensions with $z=D$. It was also realized very interestingly that the connection to conformal field theory (CFT) found in QLM is retained for general $z=D>2$. This was further generalized in [21] where field theory for general non-integer $z$ was constructed with the help of fractional derivatives. It was found that the ground state of the Lifshitz invariant scalar field continue to take the form of RK for general $z$. Furthermore, it was found that even though the usual CFT twisted operator is not available in the Lifshitz invariant field theory, due to the special property of the RK vacuum, it is possible to employ the replica trick in a path integral formalism that they developed to study the entropic properties of these theories [21]. It was found that the results are connected to CFT for more general non-integer value of $z$. Further results in holography were also obtained in [21], and a holographic proposal for boundary Lifshitz field theory in [22].
Clearly, the physics of anisotropy is a subject of its own interest that goes beyond quantum field theory. With general applicability in mind, we explore in this paper the physical effect of anisotropy in the more basic quantum mechanics. In particular we consider time anisotropy of the background space time specified by (I.1). With the dispersion relation fixed to be of the form (I.2), quantum mechanics of free particle in the anisotropic spacetime reads
$i\hbar\frac{\partial}{\partial t}\psi=H_{0}\psi,\quad H_{0}=-\frac{\hbar^{z}}{M^{z-2}}\frac{\Delta_{z}}{2m}$ (I.3)
where $H_{0}$ is the free Hamiltonian, $m$ is the mass of the particle and $M$ is a mass scale associated with the anisotropy. Here, defined by its action
$\partial_{j}^{z}e^{i{\bf k}\cdot{\bf x}}:=(ik_{j})^{z}e^{i{\bf k}\cdot{\bf x}}$ (I.4)
on the basis of plane wave functions, $\partial_{j}^{z}$ is an fractional derivative of order $z$, and $\Delta_{z}:=\partial_{j}^{z}\partial_{j}^{z}$ is the fractional Laplacian
$\triangle_{z}\,e^{i{\bf k}\cdot{\bf x}}=(-{\bf k}^{2})^{z/2}e^{i{\bf k}\cdot{\bf x}}.$ (I.5)
The free Hamiltonian (I.3) is Lifshitz scale invariant. One may also add a potential and consider the Lifshitz quantum mechanics defined by
$i\hbar\frac{\partial}{\partial t}\psi=H\psi,\quad H=-\frac{\hbar^{z}}{M^{z-2}}\frac{\Delta_{z}}{2m}+V.$ (I.6)
Note that other definitions of fractional derivatives are possible. For example, the Riesz derivative (see e.g. [24]) is defined also in terms of the Fourier transformation. However, non-analytic function $|k_{j}|$ is involved and it does not allow one to use complex contour integral theorems. For the analysis of [21], it was crucial to adopt the definition (I.4) since results such as the solution $\phi=x^{z-n}$ to the equation of motion $\partial_{x}^{z}\phi=0$ was obtained only with the help of contour deformation techniques. In this paper, we are interested in quantum mechanics that can be obtained as a low energy single particle limit of the Lifshitz QFT, therefore we consider the form (I.6) of Lifshitz quantum mechanics instead of fractional quantum mechanics defined by other types of fractional derivatives (see e.g. [24]). Note also that in general the scaling symmetry is broken except for some special form of the potential, e.g. $V\sim 1/|x|^{z}$, in which case we obtain the Lifshitz invariant quantum mechanics. In analogous to conformal quantum mechanics [23], it may be interesting to study further this kind of Lifshitz invariant quantum mechanics.
Here is a summary of the paper. In section 2, we consider the Lifshitz quantum mechanics in general. There is a belief in the literature that local conservation of probability is violated by a source term in fractional quantum mechanics. We show that this is not true for the Lifshitz quantum mechanics. In fact here the probability current is modified by anisotropy, and once the correct probability current is employed, the local conservation of probability is respected. We argue that similarly a careful identification of the probability current is needed for the other types of fractional quantum mechanics and probability is conserved in fractional quantum mechanics in general. In section 3, we consider an anisotropic Josephson Junction that is made up of an anisotropic insulating layer sandwiched between two normal superconducting materials. We show that the critical current can be significantly enhanced by anisotropy. We argue that anisotropic Josephson Junction has great potential in quantum computing.
II Lifshitz Quantum Mechanics
Consider Lifshitz quantum mechanics (I.6). This kind of quantum mechanics is also refereed to as fractional quantum mechanics, see for example [24] for some discussion. It is straightforward to show that the Lifshitz quantum mechanics (I.6) is unitary and admits the local conservation law
$\frac{\partial\rho}{\partial t}+\nabla\cdot\pmb{J}=0,$ (II.1)
where the probability density takes the usual form $\rho=\psi^{*}\psi$, while the probability current density is modified by anisotropy. Since the fractional derivative is defined using the basis of exponential functions, $\pmb{J}$ can be most easily obtained in the momentum space and we have
$\pmb{J}=-\frac{1}{2mM^{z-2}}\int\frac{d^{D}{p}\,d^{D}{q}}{(2\pi\hbar)^{D}}\,e^{i({\bf p}-{\bf q})\cdot{\bf x}/\hbar}\varphi^{*}({\bf q},t)\varphi({\bf p},t)\frac{(-{\bf q}^{2})^{z/2}-(-{\bf p}^{2})^{z/2}}{{\bf q}^{2}-{\bf p}^{2}}({\bf q}+{\bf p}),$ (II.2)
where $\varphi({\bf p},t)$ is the momentum space wave function defined by the Fourier transform
$\psi({\bf x},t)=\int\frac{d^{D}p}{(2\pi\hbar)^{D/2}}\,e^{i{\bf p}\cdot{\bf x}/\hbar}\varphi({\bf p},t).$ (II.3)
This result can be obtained from the Noether theorem associated with the phase symmetry $\psi\to\psi e^{i\theta}$ of quantum mechanics or by the direct manipulation of the wave equation (I.6). In fact, it is
$\psi^{*}\Delta_{z}\psi-\psi\Delta_{z}\psi^{*}=\int_{p}\int_{q}e^{i({\bf q}-{\bf p})\cdot{\bf x}}\varphi^{*}({\bf p})\varphi({\bf q})((-{\bf q}^{2})^{z/2}-(-{\bf p}^{2})^{z/2}),$ (II.4)
where we have denoted $\int_{p}:=\int\frac{d^{D}p}{(2\pi\hbar)^{D/2}}$. This can be identify this with a divergence term $\nabla\cdot{\bf K}$ of the form
${\bf K}=-i\int_{p}\int_{q}{\bf b}({\bf p},{\bf q})e^{i({\bf q}-{\bf p})\cdot{\bf x}}\varphi^{*}({\bf p})\varphi({\bf q})((-{\bf q}^{2})^{z/2}-(-{\bf p}^{2})^{z/2})$ (II.5)
if $\bf b$} satisfies
${\bf b}\cdot({\bf q}-{\bf p})=1.$ (II.6)
Here $\bf K$} is the current density up to a normalization factor. The solution to (II.6) is not unique in 3 spatial dimensions. For example, (II.6) can be solved by (with a suitable regulator implemented, e.g. $i\varepsilon$ in the denominator),
${\bf b}=\frac{{\bf q}+{\bf p}}{{\bf q}^{2}-{\bf p}^{2}},\quad\mbox{or}\quad{\bf b}=\frac{{\bf q}-{\bf p}}{|{\bf q}-{\bf p}|^{2}},$ (II.7)
It is easy to understand the physical meaning of the non-uniqueness. Consider any two solutions ${\bf b_{1}},{\bf b_{2}}$ of (II.6), then by definition, ${\bf b_{2}}-{\bf b_{1}}$ is perpendicular to ${\bf q}-{\bf p}$. This means
${\bf b_{2}}-{\bf b_{1}}=({\bf q}-{\bf p})\times{\bf s}$ (II.8)
for some vector $\bf s({\bf p},{\bf q})$}. This then implies that the corresponding difference ${\bf K_{2}}-{\bf K_{1}}$ is given by a curl:
${\bf K_{2}}-{\bf K_{1}}=-\nabla\times\int_{p}\int_{q}{\bf s}\;e^{i({\bf q}-{\bf p})\cdot{\bf x}}\varphi^{*}({\bf p})\varphi({\bf q})((-{\bf q}^{2})^{z/2}-(-{\bf p}^{2})^{z/2}).$ (II.9)
Therefore the different solutions of (II.6) simply give rises to current densities that are different from each other by a curl term. This does not affect the expression of current conservation. For example, for the first solution in (II.7), we obtain in coordinate space,
$\mathbf{J}=\lambda_{M}^{z-2}\frac{\hbar}{2mi}\frac{\nabla_{x}-\nabla_{x^{\prime}}}{\nabla_{x}^{2}-\nabla_{x^{\prime}}^{2}}\left[(\nabla_{x}^{2})^{z/2}-(\nabla_{x^{\prime}}^{2})^{z/2}\right]\psi^{*}(\mathbf{x}^{\prime},t)\psi(\mathbf{x},t)\Big|_{\mathbf{x}^{\prime}=\mathbf{x}},$ (II.10)
where $\lambda_{M}:=\hbar/M$ is a length scale associated with the anisotropy. We remark that this kind of ambiguity is there even for the standard quantum mechanics although it is not usually mentioned. Like the Poynting vector, additional input from the physical setting such as symmetry, boundary condition or a good understanding of the underlying dynamics, is usually needed to fix the choice of $\pmb{J}$ properly. We remark also that the local conservation of probability in fractional quantum mechanics has been discussed in [25, 26, 27] where a source term has been found, and which has led to the conclusion of a non-conservation of probability. If this is indeed the case, it posts serious doubt in the physical consistency of the fractional quantum mechanics. However as we showed above, the probability current is indeed modified in the Lifshitz quantum mechanics and local conservation of probability is observed. We believe that the probability current considered in the fractional quantum mechanics [25, 26, 27] can also be improved similarly so as to maintain probability conservation in quantum mechanics. This point is important in order to have a consistent physical interpretation of the mathematical results computed from the fractional differential equations. In the next section, we consider Lifshitz quantum mechanics for a Josephson Junction system and predict an observable effect on the macroscopic supercurrent from the conserved probability current in the insulating layer of the Josephson Junction.
III Anisotropic Josephson Junction
As an example of how anisotropy may lead to interesting quantum physics, we consider tunneling through an anisotropic region. A prominent example where the tunneling effect has a macroscopic manifestation is the Josephson effect where cooper pairs of electrons can flow between two superconductors separated by a thin insulating barrier.
Consider a SIS type Josephson Junction that is made up of two superconductors separated by a thin insulator. The two superconductors are assumed to be made of ordinary isotropic materials while the insulator is made of an anisotropic material. The latter can be achieved by material engineering such that the dispersion relation (I.2) arises and the insulator behaves effectively as an anisotropic material. In quantum mechanics, the superconductors have Landau-Ginzburg order parameters
$\psi_{i}=\sqrt{n_{i}}e^{{\rm i}\theta_{i}},\quad i=1,2$ (III.1)
that can be interpreted as the wave functions of the Cooper pairs in the two superconducting regions. See figure 1.
Figure 1: Two superconductors separated by a thin insulator
Suppose the Cooper pairs have typical energy $E$ in the regions 1, 2. For a junction with planar symmetry and that the insulator may be approximated by a potential barrier $V_{0}>E$ (see figure 2), the wave equation inside the insulator reads
$\lambda_{M}^{z-2}\frac{\hbar^{2}(\partial_{x}^{z})^{2}\psi}{2m}=(V_{0}-E)\psi,\quad-a<x<a,$ (III.2)
where $x$ is the direction perpendicular to the plane of the layers. Using the definition (I.4) of the fractional derivative, we obtain the general solution
$\psi=C_{1}\cosh\frac{x}{\zeta}+C_{2}\sinh\frac{x}{\zeta},$ (III.3)
where $C_{1},C_{2}$ are arbitrary constants and
$\zeta:=\zeta_{0}\left(\frac{\zeta_{0}}{\lambda_{M}}\right)^{2/z-1},\quad\mbox{with}\;\;\zeta_{0}:=\frac{\hbar}{\sqrt{2m(V_{0}-E)}}.$ (III.4)
Here $\zeta$ is the tunneling penetration length and $\zeta_{0}$ is the tunneling penetration length in the standard Schrödinger $(z=2)$ case. Note that the penetration length is modified by anisotropy through the presence of the anisotropy length scale $\lambda_{M}$ and anisotropy index $z$. Next let us determine the tunneling current.
Figure 2: Tunneling through a potential
Note that since we have assumed the Junction to be uniform in the transverse directions, the system is effectively one dimension and so the current is uniquely defined. From (II.10) we obtain
$J=\frac{z\zeta}{2\zeta_{0}}\cdot\frac{\hbar}{m\zeta_{0}}{\rm Im}(C_{1}^{*}C_{2}).$ (III.5)
The coefficients $C_{1},C_{2}$ can be determined by imposing the boundary conditions that the wavefunction matches up with the Landau-Ginzburg order parameters (III.1) of the superconductors at $x=\pm a$:
$\psi(-a)=\sqrt{n_{1}}e^{{\rm i}\theta_{1}},\quad\psi(a)=\sqrt{n_{2}}e^{{\rm i}\theta_{2}}.$ (III.6)
Taking into account the Cooper pair has charge $q=2e$, we obtain the first Josephson equation for the tunneling current ${\mathrm{J}}:=qJ$:
${\mathrm{J}}={\mathrm{J}}_{c}\sin\varphi,\quad\varphi=\theta_{2}-\theta_{1}.$ (III.7)
Here the critical current ${\mathrm{J}}_{c}$ is given by
${\mathrm{J}}_{c}=\gamma{\mathrm{J}}_{c0}$ (III.8)
with
${\mathrm{J}}_{c0}=\frac{q\hbar\sqrt{n_{1}n_{2}}}{m\zeta_{0}}\frac{1}{\sinh 2a/\zeta_{0}},\quad\gamma=\frac{z\zeta}{2\zeta_{0}}\frac{\sinh 2a/\zeta_{0}}{\sinh 2a/\zeta},$ (III.9)
where ${\mathrm{J}}_{c0}$ is the critical current in the isotropic case and $\gamma$ is an modification factor due to anisotropy. If we consider the application of a voltage $V$ across the Josephson Junction, then the phase becomes time dependent and we have the second Josephson equation,
$\frac{\partial\varphi}{\partial t}=\frac{q}{\hbar}V.$ (III.10)
Note that this result is unmodified by anisotropy. The derivation is the same as usual. In fact, due to gauge invariance, the wavefunction in the superconductor is modified to have a phase factor
$\psi\to\psi\exp\left(\frac{iq}{\hbar c}\int{\bf A}\cdot d{\bf x}\right)$ (III.11)
when an electromagnetic field is introduced. This results in a modification to the order parameter
$\theta_{2}-\theta_{1}=(\theta_{2}-\theta_{1})_{0}+\frac{q}{\hbar c}\int_{1}^{2}{\bf A}\cdot d{\bf x},$ (III.12)
where the subscript “0” denotes the quantity before the electric field is introduced. This leads immediately to (III.10).
Note that in the usual isotropic case, the critical current ${\mathrm{J}}_{c0}$ is exponentially suppressed by a factor of $e^{-2a/\zeta}$ since the Cooper pairs cannot tunnel deep enough into the insulating region. However this may change in the anisotropic situation. In fact, the Josephson current can be enhanced ($\gamma>1$) if the penetration length is made bigger ($\zeta>\zeta_{0}$). This happens if $z<2$ and $\lambda_{M}<\zeta_{0}$ or $z>2$ and $\lambda_{M}>\zeta_{0}$. Thus an anisotropic material has the potential of enhancing the working of Josephson Junction. In particular, if the anisotropic insulating material can be engineered such that the penetration length $\zeta$ get to become of the same order as $a$, then tunneling can become much more effective and ${\mathrm{J}}_{c}$ may be significantly enhanced. As illustration, let us consider an isotropic sample with $\zeta_{0}/2a=1/10$, in this case the critical current is suppressed with a factor of $\sim e^{-10}\sim 10^{-5}$. Introducing anisotropy, for example with $z=0.8$ and $\zeta_{0}/\lambda_{M}=10$, we can have a huge amplification factor $\gamma\sim 10^{5}$; or with $z=5$ and $\zeta_{0}/\lambda_{M}=1/10$, we have $\gamma\sim 10^{4}$. We remark that the dispersion relation (I.2) is acausal for $z<1$ due to the existence of superluminal modes [28] and the violation of the null energy conditions of the holographic dual [29]. Therefore a junction with $z\geq 1$ may be more physical. It is interesting to verify this prediction of enhanced tunneling experimentally.
The amplification of critical current due to time anisotropy may be useful for improving the performance of Josephson Junction in quantum computing. Josephson Junction also found useful application in digital circuits as switching elements due to it’s high switching speed and low power dissipation [30]. Some characteristic time constants in Josephson Junction depends on the critical current, an anisotropic Junction can have time constants that are more useful for the switching purpose. Given the wide variety of applications of Josephson Junction [31], anisotropic Josephson Junction appears to be an interesting topic that deserves further study.
Acknowledgments
CSC would like to thank Harold Steinacker for discussions and the physics department of the University of Vienna for hospitality where part of the work was done. The support of this work by NCTS and the grant 113-2112-M-007-039-MY3 of the National Science and Technology Council of Taiwan is gratefully acknowledged.